# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ 9 Conclusions and Future Directions
Apart from the above-stated research mentioned in this survey, there is still much work to be done in this field. Next, our goal is to share some insights on future directions, highlighting what is worth exploring to further this field.
Domain Adaptation. The training process of conventional deep learning based models is generally based on the i.i.d. assumption that the training and testing samples have identical and independent distribution (Schölkopf et al., 2007). Generally, as data-driven techniques, conventional deep learning based methods rely heavily on the diversity of training data to adapt to different scenarios. However, collecting and annotating enough RS images of all possible domains for model training is expensive, time-consuming, and even prohibitively impossible, especially in the military field (Wang and Deng, 2018; Wang et al., 2021a). In real-world RS applications, oriented object detectors may need to process various images collected at different times (day or night, summer or winter), in different sensors (e.g., SAR, optical, LIDAR, infrared), or under different conditions (weather, illuminance, camera pose, image quality), resulting in the domain distribution gaps between training images and testing images and degrading the performance at test time. To this end, it is desirable to investigate useful and efficient domain adaptation theories and techniques that can guide model design and enhance the generalization capability of the detectors. Recently, a large amount of valuable and instructive methods for domain adaptation have achieved many inspiring results on visual tasks, including discrepancy-based methods (Yosinski et al., 2014), adversarial-based methods (Goodfellow et al., 2014; Tzeng et al., 2017), and reconstruction-based methods (Ghifary et al., 2015, 2016).
Multimodal Information Fusion. Multimodal data has become easy to access in both military and civil fields with the rapid development of a variety of RS technologies, such as Global Positioning System (GPS), Inertial Measurement Unit (IMU), UAV, satellite, and various sensors. In recent years, the most popular multimodal information fusion methods are based on attention mechanisms (Nam et al., 2017) and bilinear pooling (Ben-younes et al., 2017), which have provided remarkable improvements for applications related to image captioning (Vinyals et al., 2015), text-to-image generation (Zhu et al., 2019), and visual question answering (Wang et al., 2018). Although multimodal information fusion methods are attracting growing attention (Zhang et al., 2020a), there are still several challenging problems in the fields of RS and oriented object detection that are far from being solved. The first problem is how to integrate various multimodal RS information, including the poses of sensors, multispectral data, and point clouds, to improve detection performance, since these data are drastically different from each other naturally. Furthermore, multimodal information fusion technologies are important in extending oriented object detection to other RS applications, including instance segmentation (Chen et al., 2019) and object tracking (Wen et al., 2021). Finally, transferring well-trained detectors to other modalities is also a challenging task.
Lightweight Methods. The demand for deploying a real-time object detector on resource-constrained mobile devices has grown rapidly. Therefore, achieving a better trade-off between accuracy and efficiency for detectors is of significant importance. To this end, various lightweight architectures have been proposed to reduce computational complexity and spatial complexity without harming accuracy. Several works try to redesign lightweight network architectures based on hand-crafted technologies (Jiang et al., 2022) or Neural Architecture Search (Xiong et al., 2021). Another branch of work presents various compression schemes for compressing networks, including parameter pruning (Hanson and Pratt, 1988; Han et al., 2015), quantization (Song et al., 2016), and knowledge distillation (Hinton et al., 2015). Excellent lightweight generic object detection methods (Xiong et al., 2021; Jiang et al., 2022; Mehta and Rastegari, 2021) depend on the ingenious designs of lightweight feature extraction networks and the efficient information transmission mechanism. Such practices provide an enormous reference value for oriented object detection. In addition, different lightweight methods may be synergistic and complementary to each other. Thus, reasonable design and combination can achieve outstanding performance. The development of lightweight methods is conducive to the promotion and real-world application of oriented object detection in the RS field, which drives the requirement for further research.
@@
The -based representation adopts a vector in the format of to define an OBB. The present approaches can be classified into two types according to the definition of the angle , including (which follows the OpenCV protocol) and the long edge definition. As shown in Figure 9, the former defines as the acute (or right) angle between the OBB and -axis, leading to . Note that the width is defined as the side of the acute angle and can be shorter than the height , which is shown at the top of Figure 9. To tackle this issue, the long edge definition is proposed by setting as the angle between the long edge of the OBB and -axis. Therefore, the angular range is (Ding et al., 2019; Han et al., 2021a) or (Han et al., 2022), which are shown in Figure 9 and Figure 9, respectively. As shown in the bottom of Figure 9, the parameters of the same OBB have significant differences in different OBB representations.
@@
where denotes the norm and the smooth loss (Girshick, 2015) is widely adopted. Due to the PoA (Qian et al., 2021, 2022; Yang et al., 2021c, 2022b), the OBB regression will encounter the following challenges.
(1) Inconsistency between Metric and Loss. Although the majority of detectors adopt the smooth L1 loss as the objective function of regression, the most commonly used metric for localization is RIoU. Therefore, there is an inconsistency between the loss function and the evaluation metric. This implies that an optimum choice for the regression task may not guarantee a high localization accuracy in terms of RIoU. What’s more, a good regression loss function should take into account the central point distance, aspect ratio, and overlap area, which has been demonstrated to be effective in horizontal object detection (Rezatofighi et al., 2019; Zheng et al., 2020). However, the aspect ratio and the overlap area can be disregarded by the smooth L1 loss easily.
We illustrate the inconsistency between metric and loss in Figure 10. As shown in Figure 10, the top and the bottom rows have different angle differences, while the aspect ratio of the OBBs on the left is different from those on the right. Meanwhile, the center points, width, and height of the four cases are the same. The orange area denotes the IoU between OBBs. Note that the regression loss is sensitive to angle variances but remains unchanged for different aspect ratios. Specifically, when the aspect ratio varies, the union of two OBBs will change but the intersection is constant, causing the change of RIoU. The same conclusion can be drawn from Figure 10, which shows the variation curves of the RIoU and smooth L1 loss w.r.t aspect ratio under different angle differences. Note that the RIoU changes drastically but the smooth L1 loss remains constant. Furthermore, Figure 10 shows the variation curve of RIoU and smooth L1 loss with respect to the angle under different aspect ratios. In the neighborhood of 0, both losses are consistent in monotonicity but not in convexity. The RIoU changes more intensely than the smooth L1 loss when the angle difference is close to zero.
@@
Because of the PoA problem (Yang et al., 2021c, d, 2022; Qian et al., 2021, 2022), the smooth L1 loss suffers from the problem of angular boundary discontinuity, which is illustrated in Figure 11. Specifically, a small angle difference may cause a large loss change when the angular value approaches the angular boundary range. Figure 11 illustrates an ideal OBB representation, where the predicted and GT OBB only differ slightly in terms of the angle and center point. For OBBs with OpenCV definition, the angular value must be an acute or right angle, i.e., , as shown in Figure 11. As a result, the angle difference between the two OBBs increases sharply as the angular value is close to 0 or . In addition, the width of the predicted OBB is the short edge, whereas the width of the GT OBB is the long edge, causing a significant regression loss of width and height. For OBBs under long edge definition with the angular range of , the angular boundary discontinuity leads to a significant angle difference, i.e., , as shown in Figure 11. This problem will also occur in long edge definition with the angular range of when the angular value is close to or .
For square-like objects, including storage-tank and roundabouts, the long edge definition will encounter a so-called square-like problem due to the difference of angle parameters (Yang et al., 2021c, d, 2022). As shown in Figure 12, when the aspect ratio is close to 1 but the length and width of the predicted OBB are opposite to that of GT, the corresponding angle will differ by about , leading to a large regression loss even if the RIoU is about 1.
@@ Appendix B Quadrilateral Representation
The quadrilateral representation denotes an OBB as a vector , where indicates the image coordinates of the vertex arranged in a clockwise order (Xu et al., 2021a). This representation method can compactly enclose oriented objects with large deformation and has been widely adopted to annotate objects in large-scale RS datasets, including DOTA (Xia et al., 2018; Ding et al., 2022), and HRSC2016 (Liu et al., 2016b). Significantly, the top-left vertex relative to the object orientation is chosen as the starting point ), as shown in Figure 13(a).
@@
Generally, the anchor box selects the top-left vertex in the image as the starting point. To ensure consistency, the leftmost vertexes of the predicted OBB and the corresponding GT OBB are chosen as the starting point, as shown in Figure 13(b). However, the inappropriate vertex sorting may cause inconsistencies between the vertex sequences of the anchor and the GT OBB, which is known as the vertexes sorting problem or the corners sorting problem (Qian et al., 2021, 2022; Xu et al., 2021a). Figure 14 shows a case of the problem. The anchor and the GT OBB are shown in blue and green respectively, and the dashed line and the solid line denote the actual and ideal vertexes matching during regression. In the ideal setting, the vertexes matching from the anchor to the GT is: . However, in the actual regression, the vertexes matching is: . Such inconsistency causes a large regression loss, confusing the network during the training process. Hence, it is critical to determine the sequence of vertexes in advance to stabilize the training process.
@@ Appendix C Datasets
In Table 2 of our main paper, we review a series of benchmarks regarding oriented object detection. However, space constraints prevent us from presenting all of them in detail. In this section, further details regarding the datasets mentioned in Section 7.1 are presented.
@@
SZTAKI-INRIA (Benedek et al., 2012) contains 665 buildings in 9 multi-sensor aerial or satellite images taken from different cities. Due to the small capacity, this dataset is used to evaluate traditional object detection algorithms.
@@
3K vehicle (Liu and Mattyus, 2015) is created for vehicle detection, comprising 20 images and 14,235 vehicles. The images have a resolution of and are captured by a DLR camera system at a height of 1,000m above the ground. Therefore, the ground sample distance (GSD) is approximately 13 cm, leading to smaller scale variations. Besides, the images have a similar background. Hence, this dataset is excluded from the evaluation of algorithms on complicated scenes.
@@
VEDAI (Razakarivony and Jurie, 2016) is also proposed for vehicle detection, containing more categories and a wider variety of backgrounds, e.g., fields, grass, mountains, urban area, etc, making the detection more complicated. It comprises 1,210 images with a resolution of . The images are cropped from Very-High-Resolution (VHR) satellite images with a GSD of 12.5cm. However, the dataset only consists of 3,640 instances, because the images with too many dense vehicles are excluded. It is worth mentioning that each image has four color channels, including three visible channels and one 8-bit near-infrared channel.
@@
UCAS-AOD (Zhu et al., 2015) contains 7,482 planes in 1,000 images, 7,114 cars in 510 images, and 910 negative images. All images in this dataset are cropped from Google Earth aerial images. Especially, the instances are carefully selected to ensure their orientations are distributed evenly.
@@
DOTA (Xia et al., 2018; Ding et al., 2022). Figure 15 shows the number of instances for each category in training and validation subsets of DOTA-V1.0, V1.5, and V2.0. Note that the distributions of different categories are severely imbalanced. The instances of small-vehicle and ship have a large quantity, while nearly half of the other categories have quantities of less than 1,000, including plane, baseball diamond, ground track field, basketball court, soccer ball field, roundabout, helicopter, container crane, airport and helipad. The severe category imbalance makes the model seriously over-fitting to the many-shot categories but under-fitting to the low-shot categories (Gupta et al., 2019; Cui et al., 2019; Wang et al., 2021b). Figure 16 further summarizes the size and ratio distributions for each category in three versions of DOTA, respectively. As shown in Figure 16(a), the minimum size is orders of magnitude lower than the maximum size in each category. Moreover, there is also a large range of size differences between categories. Figure 16(b) indicates that the aspect ratios of different categories vary greatly. Furthermore, some categories have an extremely large aspect ratio, such as bridge, harbor, and airport. Up to now, DOTA has been the most challenging dataset for oriented object detection, due to its tremendous object instances, large aspect ratio, significant size variance, and complicated aerial scenes. All of these characteristics contribute to DOTA as the de facto benchmark for evaluating the efficacy of oriented object detectors in previous years.

### v5

@@ 9 Conclusions and Future Directions
Despite the rising prominence of artificial intelligence, deep learning has rapidly advanced the development of oriented object detection methods. However, due to the complex and ever-changing real-world scenarios, its performance still faces limitations, hindering robust and reliable practical applications. Over the years, a considerable amount of research effort has been dedicated to tackling the challenges of feature misalignment, spatial misalignment, OBB regression, as well as common issues (e.g., complex background, scale variations, large aspect ratio, and lack of annotated samples). These efforts have led to marked improvements in detection performance. Nevertheless, it is imperative to acknowledge that a considerable gap persists between the current detection capabilities and the demands of practical applications. More critically, several issues remain insufficiently addressed, posing significant barriers to further advancements in oriented object detection technology.
Low detection efficiency. Detection efficiency stands as a pivotal factor in the real-world applications of detectors. Current state-of-the-art oriented object detection models are designed to be exceptional complexity to achieve superior detection accuracy. Nevertheless, their intricate network architectures markedly impede detection efficiency, rendering them unsuitable for real-time applications.
Imbalance in datasets. Modern datasets focus on general scenarios and contain images from a variety of different environments and contexts. While this diversity helps models learn a wider range of features, it may also lead to poor performance in specific scenarios (such as snow, fog, occlusion). Additionally, common scenes or objects may constitute the majority of the datasets, while rare or special scenes or objects may be scarce. This data imbalance issue may cause the model to develop biases towards certain scenes or objects during training, resulting in a degradation of performance.
Detection in single-model image. The current research community is dedicated to exploring and developing oriented object detection methods in single-model image. However, these methods are inherently constrained by their reliance on a solitary information source and the lack of contextual cues, exhibiting a heightened vulnerability to various interferences, including variations in lighting conditions, occlusions, and shadows.
Given that the aforementioned issues have not yet been explored and studied in oriented object detection, we further share some insights on future directions.
Lightweight methods. The demand for real-time object detectors on resource-limited mobile devices is on the rise, necessitating innovative solutions to overcome hardware constraints. Thus, light-weight oriented object detection architectures are required to fulfill the requirement of mobile and embedded applications. To promote the application of oriented object detection in real-world scenarios, a feasible approach is to adopt meticulously designed efficient network architectures or leverage Neural Architecture Search [178] to discover optimal architectures. These light-weight network architectures enhance the efficiency and accuracy of feature extraction, while concurrently reducing model parameter count and computational burden. For example, several lightweight network structures, such as MobileNet [64] and ShuffleNet [232], have gained widespread adoption in object detection tasks Another feasible method is model compression to develop highly competitive, compact, and rapid detection models, including parameter pruning [58, 59, 39, 212], quantization [140, 179, 24], and knowledge distillation [62, 226, 154]. These compression techniques have demonstrated remarkable effectiveness in bolstering generalization capabilities and mitigating underfitting during the training of efficient object detection models. These compression techniques have demonstrated remarkable effectiveness in improving generalization capabilities and mitigating underfitting during the training of object detection models.
Mission-specific datasets. In light of the prevailing imbalance in categories and scenarios within current datasets, coupled with the emerging trend of research on multi-modal large-scale models, we will delineate the future direction for datasets collection from three aspects: scenario-specific datasets, multi-modal datasets, and large-scale datasets.
Scenario-specific datasets can provide more refined and accurate data tailored to specific scenarios (e.g., severe weather conditions, or rare scenarios), empowering models to achieve superior performance within those specific scenarios. Moreover, these datasets can effectively tackle the challenge of high-quality data scarcity in specialized scenarios, thereby enhancing models’ generalization capabilities and promoting the practical application of oriented object detection in specialized fields.
Compared to single-model-based oriented object detection datasets, multi-modal ones integrate diverse data types that complement each other, providing a wealth of data resources essential for tackling intricate problems. By leveraging the connections and relationships among various data types, and fusing information from multiple modalities, the accuracy and performance of models can be substantially enhanced, leading to more comprehensive and accurate analysis results. An interesting attempt is the DroneVehicle [136], in which the two modalities—RGB and infrared—offer complementary information across different lighting conditions. RGB images offer rich color information, while infrared images excel in low-light conditions and are not affected by darkness. In the future, with the continuous development of more efficient and accurate multi-modal technologies [12, 133, 89, 88], multi-modal datasets will play a pivotal role in a wide array of fields, driving continuous innovation in oriented object detection technology.
Owing to the exceptional representation and generalization capacities, large models have emerged as a focal point of current research [151, 51, 85, 69]. Relying on extensive data samples, large models can discerning intricate features and underlying patterns within the data, thereby enhancing the both the accuracy and generalization prowess. As the technology pertaining to large models continues to advance, their ability to handle complex scenarios and multi-modal data will be further enhanced, which will bring more possibilities to oriented object detection. Consequently, there is a pressing need to develop high-quality, large-scale datasets tailored for oriented object detection. These datasets will not only provide a rich repository of samples but also serve as a robust validation platform to facilitate the training and evaluation of large models in this domain.
Multi-modal large models. Multi-modal large models, as a critical avenue for the progression of artificial intelligence towards Artificial General Intelligence (AGI), have emerged as a focal point of current research [88, 69]. . However, existing methods for oriented object detection predominantly concentrates on extracting valuable information from single-model images, neglecting the guidance provided by multi-modal data. On the other hand, the scarcity of high-quality multi-modal data is a fundamental bottleneck that impedes the advancement of multi-modal large models within the domain of oriented object detection. Looking ahead, a prominent direction for future research entails the exploration and integration of multi-modal data. This encompasses the amalgamation of text and visual large models, the incorporation of GPS, IMU, and remote sensing imagery, as well as the fusion of diverse sensors. By harnessing the synergies among these various data modalities, we can unlock new potentials for enhancing the accuracy and robustness of oriented object detection methods. As technology continues to evolve, driven by the proliferation of data, and the expansion of application scenarios, multi-modal large models are expected to significantly elevate the performance of oriented object detection.
@@
The -based representation adopts a vector in the format of to define an OBB. The present approaches can be classified into two types according to the definition of the angle , including OpenCV definition (which follows the OpenCV protocol) and the long edge definition. As shown in Fig. 16a, the former defines as the acute (or right) angle between the OBB and -axis, leading to . Note that the width is defined as the side of the acute angle and can be shorter than the height , which is shown at the top of Fig. 16a. To tackle this issue, the long edge definition is proposed by setting as the angle between the long edge of the OBB and -axis. Therefore, the angular range is [29, 48] or [47], which are shown in Fig. 16b and Fig. 16c, respectively. As shown in the bottom of Fig. 16, the parameters of the same OBB have significant differences in different OBB representations.
@@
where denotes the norm and the smooth loss [37] is widely adopted. Due to the PoA [125, 126, 198, 202], the OBB regression will encounter the following challenges.
(1) Inconsistency between Metric and Loss. Although the majority of detectors adopt the smooth L1 loss as the objective function of regression, the most commonly used metric for localization is RIoU. Therefore, there is an inconsistency between the loss function and the evaluation metric. This implies that an optimum choice for the regression task may not guarantee a high localization accuracy in terms of RIoU. What’s more, a good regression loss function should take into account the central point distance, aspect ratio, and overlap area, which has been demonstrated to be effective in horizontal object detection [134, 220]. However, the aspect ratio and the overlap area can be disregarded by the smooth L1 loss easily.
We illustrate the inconsistency between metric and loss in Fig. 17. As shown in Fig. 17a, the top and the bottom rows have different angle differences, while the aspect ratio of the OBBs on the left is different from those on the right. Meanwhile, the center points, width, and height of the four cases are the same. The orange area denotes the IoU between OBBs. Note that the regression loss is sensitive to angle variances but remains unchanged for different aspect ratios. Specifically, when the aspect ratio varies, the union of two OBBs will change but the intersection is constant, causing the change of RIoU. The same conclusion can be drawn from Fig. 17b, which shows the variation curves of the RIoU and smooth L1 loss w.r.t aspect ratio under different angle differences. Note that the RIoU changes drastically but the smooth L1 loss remains constant. Furthermore, Fig. 17c shows the variation curve of RIoU and smooth L1 loss w.r.t the angle under different aspect ratios. In the neighborhood of 0, both losses are consistent in monotonicity but not in convexity. The RIoU changes more intensely than the smooth L1 loss when the angle difference is close to zero.
@@
Because of the PoA problem [200, 203, 125, 126], the smooth L1 loss suffers from the problem of angular boundary discontinuity, which is illustrated in Fig. 18. Specifically, a small angle difference may cause a large loss change when the angular value approaches the angular boundary range. Fig. 18a illustrates an ideal OBB representation, where the predicted and GT OBB only differ slightly in terms of the angle and center point. For OBBs with OpenCV definition, the angular value must be an acute or right angle, i.e., , as shown in Fig. 18b. As a result, the angle difference between the two OBBs increases sharply as the angular value is close to 0 or . In addition, the width of the predicted OBB is the short edge, whereas the width of the GT OBB is the long edge, causing a significant regression loss of width and height. For OBBs under long edge definition with the angular range of , the angular boundary discontinuity leads to a significant angle difference, i.e., , as shown in Fig. 18c. This problem will also occur in long edge definition with the angular range of when the angular value is close to or .
For square-like objects, including storage-tank and roundabouts, the long edge definition will encounter a so-called square-like problem due to the difference of angle parameters [198, 200, 203]. As shown in Fig. 19, when the aspect ratio is close to 1 but the length and width of the predicted OBB are opposite to that of GT, the corresponding angle will differ by about , leading to a large regression loss even if the RIoU is about 1.
@@ Appendix B Quadrilateral Representation
The quadrilateral representation denotes an OBB as a vector , where indicates the image coordinates of the vertex arranged in a clockwise order [176]. This representation method can compactly enclose oriented objects with large deformation and has been widely adopted to annotate objects in large-scale RS datasets, including DOTA [172, 30], and HRSC2016 [100]. Significantly, the top-left vertex relative to the object orientation is chosen as the starting point ), as shown in Fig. 20a.
@@
Generally, the anchor box selects the top-left vertex in the image as the starting point. To ensure consistency, the leftmost vertexes of the predicted OBB and the corresponding GT OBB are chosen as the starting point, as shown in Fig. 20b. However, the inappropriate vertex sorting may cause inconsistencies between the vertex sequences of the anchor and the GT OBB, which is known as the vertexes sorting problem or the corners sorting problem [125, 126, 176]. Fig. 21 shows a case of the problem. The anchor and the GT OBB are shown in blue and green respectively, and the dashed line and the solid line denote the actual and ideal vertexes matching during regression. In the ideal setting, the vertexes matching from the anchor to the GT is: . However, in the actual regression, the vertexes matching is: . Such inconsistency causes a large regression loss, confusing the network during the training process. Hence, it is critical to determine the sequence of vertexes in advance to stabilize the training process.
@@ Appendix C Datasets
In Tab. 3 of our main paper, we review a series of benchmarks regarding oriented object detection. However, space constraints prevent us from presenting all of them in detail. In this section, further details regarding the datasets mentioned in Sec. 7.2 are presented.
@@
SZTAKI-INRIA [1] contains 665 buildings in 9 multi-sensor aerial or satellite images taken from different cities. Due to the small capacity, this dataset is used to evaluate traditional object detection algorithms.
@@
3K vehicle [93] is created for vehicle detection, comprising 20 images and 14,235 vehicles. The images have a resolution of and are captured by a DLR camera system at a height of 1,000m above the ground. Therefore, the ground sample distance (GSD) is approximately 13 cm, leading to smaller scale variations. Besides, the images have a similar background. Hence, this dataset is excluded from the evaluation of algorithms on complicated scenes.
@@
VEDAI [132] is also proposed for vehicle detection, containing more categories and a wider variety of backgrounds, e.g., fields, grass, mountains, urban area, etc, making the detection more complicated. It comprises 1,210 images with a resolution of . The images are cropped from Very-High-Resolution (VHR) satellite images with a GSD of 12.5cm. However, the dataset only consists of 3,640 instances, because the images with too many dense vehicles are excluded. It is worth mentioning that each image has four color channels, including three visible channels and one 8-bit near-infrared channel.
@@
UCAS-AOD [204] contains 7,482 planes in 1,000 images, 7,114 cars in 510 images, and 910 negative images. All images in this dataset are cropped from Google Earth aerial images. Especially, the instances are carefully selected to ensure their orientations are distributed evenly.
@@
DOTA [172, 30]. Fig. 22 shows the number of instances for each category in training and validation subsets of DOTA-V1.0, V1.5, and V2.0. Note that the distributions of different categories are severely imbalanced. The instances of small-vehicle and ship have a large quantity, while nearly half of the other categories have quantities of less than 1,000, including plane, baseball diamond, ground track field, basketball court, soccer ball field, roundabout, helicopter, container crane, airport and helipad. The severe category imbalance makes the model seriously over-fitting to the many-shot categories but under-fitting to the low-shot categories [36, 10, 171]. Fig. 23 further summarizes the size and ratio distributions for each category in three versions of DOTA, respectively. As shown in Fig. 23a, the minimum size is orders of magnitude lower than the maximum size in each category. Moreover, there is also a large range of size differences between categories. Fig. 23b indicates that the aspect ratios of different categories vary greatly. Furthermore, some categories have an extremely large aspect ratio, such as bridge, harbor, and airport. Up to now, DOTA has been the most challenging dataset for oriented object detection, due to its tremendous object instances, large aspect ratio, significant size variance, and complicated aerial scenes. All of these characteristics contribute to DOTA as the de facto benchmark for evaluating the efficacy of oriented object detectors in previous years.